# Graph-Based Bidirectional Transformer Decision Threshold Adjustment Algorithm for Class-Imbalanced Molecular Data

### v1

@@ Abstract
Data sets with imbalanced class sizes, often where one class size is much smaller than that of others, occur extremely often in various applications, including those with biological foundations, such as drug discovery and disease diagnosis. Thus, it is extremely important to be able to identify data elements of classes of various sizes, as a failure to detect can result in heavy costs. However, many data classification algorithms do not perform well on imbalanced data sets as they often fail to detect elements belonging to underrepresented classes. In this paper, we propose the BTDT-MBO algorithm, incorporating Merriman-Bence-Osher (MBO) techniques and a bidirectional transformer, as well as distance correlation and decision threshold adjustments, for data classification problems on highly imbalanced molecular data sets, where the sizes of the classes vary greatly. The proposed method not only integrates adjustments in the classification threshold for the MBO algorithm in order to help deal with the class imbalance, but also uses a bidirectional transformer model based on an attention mechanism for self-supervised learning. Additionally, the method implements distance correlation as a weight function for the similarity graph-based framework on which the adjusted MBO algorithm operates. The proposed model is validated using six molecular data sets, and we also provide a thorough comparison to other competing algorithms. The computational experiments show that the proposed method performs better than competing techniques even when the class imbalance ratio is very high.
Keywords: Imbalanced data, molecular data, transformer, graph-based, data classification, distance correlation.
@@ 1 Introduction
In the data classification task, imbalanced data is characterized by an underrepresentation of one class (the “ minority” class) in the data compared to the other class or classes (the “ majority” class(es)). As a result, data classification algorithms trained on imbalanced data tend to become biased toward the majority class(es) and fail to correctly classify data points from the minority class, which is frequently the class of most interest [12, 27]. This scenario is quite common in classification tasks, and it can be especially problematic if the minority class represents an important or critical outcome, such as in drug discovery, disease diagnosis, fraud detection, energy management, and other applications involving rare events [27]. For example, drug discovery is one of the most challenging tasks in science. In drug screening, the number of inactive molecules is typically thousands of times larger than that of active molecules. A successful drug must be highly potent, be free of serious side effects or off-target binding, and exhibit minimal or no toxicity [12]. As a result, the majority of drug candidates are either inactive or unqualified, leading to highly imbalanced data sets [39, 26]. More specifically, imbalanced data occur during virtual screening: identifying potential drug candidates from large compound libraries. Effective handling of imbalanced data can improve the identification of active compounds [37, 61, 58, 48]. Another situation in which we deal with imbalanced data is Quantitative Structure-Activity Relationship (QSAR) modeling, which predicts the activity of compounds based on their chemical structure [6, 25]. Balancing the data set can lead to more reliable predictions [53].
@@
Therefore, addressing imbalanced data is crucial for effective machine learning in many applications such as drug design and discovery. By employing a combination of data-level and algorithm-level strategies, and focusing on appropriate evaluation metrics, researchers can build models that better identify rare but valuable compounds, leading to more successful outcomes in drug discovery and development.
@@
Techniques to deal with imbalanced data can be categorized into four groups [27, 34, 56, 20]: data preprocessing or resampling methods[14, 22], model-level techniques [27, 16], ensemble algorithms [19], and cost-sensitive methods [10]. Strategies in the first group include resampling methods, which attempt to rebalance the class distribution in the training set using sampling techniques, and feature selection methods, which are carried out in the feature space rather than the sample space [27]. Examples of resampling procedures include oversampling [5, 49], undersampling [38, 2] and hybrid-sampling algorithms [14, 54, 7]. Respectively, these procedures attempt to rebalance the sample space by creating new samples from the minority class, discarding samples from the majority class, or a combination of both. Model-level techniques attempt to make adjustments to steps of the algorithm itself so that there is less bias towards classes of larger size. Ensemble methods use the output of multiple classifiers to make predictions, which can improve the performance of individual classifiers [27, 40]. Cost-sensitive methods instead assign a larger misclassification cost to classes of smaller size [10].
@@
Some cost-sensitive approaches to imbalanced data classification involve thresholding, which is applied in a postprocessing step [52]. In particular, thresholding algorithms are cost-sensitive techniques which assign a probability to each of the test elements. For a threshold of 0.5, which is the default probability threshold value in most classifiers [62], any test element with a probability of at least 0.5 will be assigned to one class; otherwise, it will be assigned to another class. In the case of highly imbalanced data classification, the threshold of 0.5 may not be optimal; thus, threshold algorithms attempt to find a more optimal threshold boundary for the machine learning task. Usually, a balanced accuracy metric is used; some examples include the ROC curve (receiver operating characteristic curve) [15], the G-Mean [33], F1-score [62], the Matthews correlation coefficient [13] and the balanced accuracy [37]. In these cases, the goal is to maximize the balanced accuracy metric.
@@
However, most of the cost-sensitive techniques for imbalanced data classification that search for the optimal threshold can be computationally extensive because retraining of the classifier is often required. Inspired by the work of Esposito et al. [12], in this paper, we detail an algorithm for imbalanced data classification which does not require any retraining in its procedure. In order to motivate the proposed procedure, we turn to the authors’ previous work [28], which demonstrated the success of their BT-MBO method for molecular data classification with very low amounts of labeled data; refer to [9] for a survey of machine learning methods for small data challenges in molecular science. The BT-MBO algorithm, which admits 2D molecular data in simplified molecular input line entry specification (SMILES) [60] format, is composed of a bidirectional transformer [8] and a graph-based Merriman-Bence-Oscher (MBO) algorithm [23, 47] in series, utilizing the transformer model to generate molecular fingerprints which then serve as features for the MBO algorithm. The MBO method used in the BT-MBO model [28] generates a probability distribution over all classes for each data point, and it subsequently classifies each data element by choosing the class for which that element has the highest probability. However, in the case of imbalanced data, this procedure may cause the model to misclassify data elements in the minority class, since the probability distribution may be biased toward the majority classes. Thus, in the present work, we propose the BTDT-MBO method (bidirectional transformer with MBO techniques using a varying decision threshold), which is a novel adaptation of the BT-MBO technique for use on highly imbalanced molecular data.
@@
In particular, the proposed BTDT-MBO method incorporates a decision threshold adjustment in the MBO algorithm, which utilizes the probability distribution generated by the algorithm for each data point. Instead of following the typical MBO classification procedure, the BTDT-MBO algorithm tests a set of thresholds and chooses the threshold that produces the highest ROC-AUC score (area under the ROC curve) for the unlabeled points. The new thresholding step introduced in the BTDT-MBO method does not require any additional iterations of the MBO algorithm; rather, at the end of each iteration, we repeat the new thresholding step for each threshold. We refer to the modified MBO algorithm with decision threshold adjustment as the DT-MBO method. Another adaptation introduced in the BTDT-MBO method is the inclusion of the distance correlation [57, 29] as a weight function in the DT-MBO algorithm. The default weight function used in our BTDT-MBO method is a Gaussian kernel; in addition to experiments using a Gaussian weight function, we also carried out experiments using a distance correlation weight function. The methods used in the present work are further detailed in Section 2.
@@
Given our previous BT-MBO method’s predictive success for the challenging scenario of scarcely labeled data [28], we focus on the difficult task of highly imbalanced molecular data sets in the present paper. Specifically, all six data sets used for benchmarking the proposed BTDT-MBO algorithm have an imbalance ratio (IR) of 16.5 or greater. We compare our results using BTDT-MBO to those of the GHOST procedure [12] applied to four common classification algorithms: random forest, gradient boosting decision trees, extreme gradient boosting, and logistic regression. Specifically, the GHOST technique [12] is an automated procedure that can be applied to any classification algorithm and that adjusts the decision threshold of a classifier using the prediction probabilities of the training set (i.e., without needing to retrain the classifier); it achieved success in classifying imbalanced molecular data, demonstrated on data sets with varying levels of imbalance. Detailed results on the six highly imbalanced data sets, as well as discussion, are given in Section 3. The computational experiments indicate that the proposed BTDT-MBO method achieves superior results to the comparison techniques.
@@
The following is a summary of the present paper’s contributions:
  • We present a new algorithm, the BTDT-MBO method, for molecular data classification, especially designed to be very useful even in cases of highly imbalanced data sets. This procedure incorporates Merriman-Bence-Osher techniques, as well as bidirectional transformers, distance correlation, and adjustments in the classification threshold.
  • The proposed BTDT-MBO algorithm is able to perform extremely well even in the cases of highly imbalanced molecular data, where the class sizes vary greatly. This is extremely important, since highly imbalanced data sets are often found in practice.
  • The proposed method utilizes several techniques to perform well in the case of imbalanced data where class sizes vary signicantly, including adjustments in the classification threshold.
  • The performance of the proposed BTDT-MBO algorithm is demonstrated using experiments on six data sets, where the proposed method performs more accurately than the comparison methods in almost all cases. The ROC-AUC score is used as a metric.
@@ 2 Methods
In this section, we describe the methods used in the present work. In Section 2.1, we outline the details of the proposed BTDT-MBO method (bidirectional transformer with MBO techniques [23, 47] using a varying decision threshold), which adapts some of the techniques of our previous work [28]. In Section 2.2, we further discuss the distance correlation [29, 57] as a potential similarity metric for use in the BTDT-MBO method.
@@ 2.1 BTDT-MBO Algorithm
The authors’ previous work [28] demonstrated the success of their proposed BT-MBO method in molecular classification problems with low amounts of labeled data, with the BT-MBO model outperforming state-of-the-art methods on some benchmark data sets when as little as 1% of data is labeled. The present work proposes a method, called BTDT-MBO, that adapts some of the procedures of the BT-MBO algorithm for classification problems on highly imbalanced molecular data sets, where the classes of the data set vary vastly in size.
Inspired in part by the success of Esposito et al. [12] in adjusting the classification threshold for machine learning classifiers to make predictions about imbalanced molecular data, the proposed BTDT-MBO method utilizes the aspects of the BT-MBO algorithm from [28] along with a decision threshold adjustment at each iteration of the method. We denote the adjusted MBO method with varied decision threshold by DT-MBO. The proposed BTDT-MBO algorithm consists of a bidirectional transformer (created by Chen et al. [8]) and a DT-MBO algorithm in series, where the fingerprints (BT-FPs) generated by the transformer serve as input features to the DT-MBO algorithm. The BTDT-MBO algorithm is summarized in Algorithm 2, and details of the DT-MBO method are given in Algorithm 1.
@@ 2.1.1 MBO Scheme with Decision Threshold Adjustment
In this section, we establish the similarity graph-based framework our BTDT-MBO scheme uses, which adapts techniques outlined in the literature [44, 45, 23]. First, consider a graph $G=(V,E)$, where $V$ is the set of vertices representing the data set elements, and $E$ is the set of edges connecting some pairs of vertices. Furthermore, we can define a weight function $w:V\times V\to\mathbb{R}$, whose value represents the degree of similarity between two vertices of the graph. Specifically, $w(i,j)$ measures the degree of similarity of data elements $i$ and $j$ of the data set. While there are many choices for the weight function, it should be constructed such that a large value of $w(i,j)$ corresponds to a high degree of similarity between data set elements $i$ and $j$, and a small value of $w(i,j)$ corresponds to a low degree of similarity between elements $i$ and $j$. Usually, the weight function depends on the particular kind of data set at hand, but a commonly used weight function [45] is the Gaussian function:
@@
where $d(i,j)$ represents the distance between data elements represented by vertices $i$ and $j$ (computed using some prespecified metric such as the Euclidean distance), and $\sigma>0$ [23, 45].
In the present work, we use the Gaussian weight function for some of our BTDT-MBO experiments, and for other experiments, we use the distance correlation [57] as a weight function. The distance correlation is defined and further discussed in Section 2.2.
Given a weight function $w$, we can further define a weight matrix $\mathbf{W}$ with elements $\mathbf{W}_{ij}=w(i,j)$ and define the degree of a vertex $i\in V$ in our graph as $d_{i}=\sum_{j\in V}w(i,j)$. If we let $\mathbf{D}$ be the diagonal matrix with diagonal elements $\{d_{i}\}$, then we can define the graph Laplacian as
@@
For outlining the Merriman-Bence-Osher (MBO) framework, originally introduced to approximate mean curvature flow in [47] and then applied using a graph-based setting to various applications such as hyperspectral imaging in works such as [44, 23, 41, 42, 46], we consider a general data classification problem with $m$ classes. Overall, the MBO algorithm ultimately aims to generate an optimal label matrix $\mathbf{U}=(\mathbf{u}_{1},\dots,\mathbf{u}_{N})^{T}$, where each row $\mathbf{u}_{i}\in\mathbb{R}^{m}$ is the probability distribution over the $m$ classes for data element $i$. For instance, the first element of $\mathbf{u}_{i}$ is the probability that data element $i$ is a member of the first class, the second element of $\mathbf{u}_{i}$ is the probability that data element $i$ is a member of the second class, and so on. Therefore, each vector $\mathbf{u}_{i}$ can be viewed as an element of the Gibbs simplex with $m$ vertices:
$\Sigma^{m}:=\{(z_{1},...,z_{m})\in[0,1]^{m}|\sum_{k=1}^{m}z_{k}=1\}\vspace{0.2cm}$ (3)
@@
The vertices of this simplex are the unit vectors $\{\mathbf{e}_{i}\}$, where the probability of belonging to class $i$ is 1, while the probability of belonging to any other class is 0. Thus, the vertices $\{\mathbf{e}_{i}\}$ correspond to data elements that belong exclusively to each of the $m$ classes. In our data classification problem with labeled and unlabeled data elements, the labeled data elements are assigned to their corresponding vertices of the Gibbs simplex.
@@
The process by which the MBO algorithm generates the optimal label matrix $\mathbf{U}$ relies on energy minimization, a time-splitting scheme, and transference to a graph-based setting as described above. A general graph-based machine learning algorithm for data classification can be formulated as minimizing the energy $\text{E}(\mathbf{U})=\text{R}(\mathbf{U})+\text{Fid}(\mathbf{U})$, where R is a regularization term that incorporates the graph weights and $\text{Fid}(\mathbf{U})$ is a fidelity term that incorporates the labeled points [45].
@@
As Garcia and coauthors demonstrated in [23], one successful choice for the regularization term R is the Ginzburg-Landau functional [4, 32], which is often used in image processing and associated applications due to its relation to the total variation functional [44, 11]. For the fidelity term $\text{Fid}(\mathbf{U})$, one can choose an $L_{2}$ fit to all elements with known class. In the continuous setting, $L_{2}$ gradient descent can then be applied to minimize the energy E, resulting in a modified Allen-Cahn equation with a forcing term. Further applying a time-splitting scheme yields a procedure that alternates between a diffusion step, which uses the heat equation with a forcing term, and a thresholding step. Finally, one can use the techniques outlined in [23, 44, 43] to transfer the procedure to a graph-based setting, in which the thresholding step is replaced by projection to the Gibbs simplex (3), followed by displacement to the nearest vertex in the Gibbs simplex. Additionally, the Laplace operator in the heat equation is replaced by the graph Laplacian (2) or some normalization of the graph Laplacian.
@@
To formulate our decision threshold technique, we are inspired by the work by Esposito et al. [12], which demonstrated success in classifying imbalanced molecular data by developing a procedure that adjusts and optimizes the decision threshold of machine learning classifiers. For binary data, the default decision threshold is usually 0.5, where the model assigns a sample to the minority class if its probability of belonging to the minority class is greater than 0.5 (this is equivalent to assigning the sample to the class for which it has the highest probability). However, machine learning classifiers are likely to misclassify minority points since the probability distribution over the two classes for a given data point may be skewed toward the majority class [36], so Esposito et al. [12] tested potential decision thresholds from 0.05 to 0.5 with a spacing of 0.05. Similarly, in the present work, we augment the MBO method by a process that optimizes the decision threshold for the displacement step, resulting in our proposed decision-threshold-MBO (DT-MBO) algorithm.
@@
Overall, we summarize the main iterative steps of the DT-MBO algorithm for the classification task used in the present work below, with more details given in Algorithm 1. The process is adapted from the MBO algorithm used in our previous work [28], with an additional decision threshold step in the displacement step. The below steps require a given labeled/unlabeled partition of a data set as well as a list of thresholds to test.
@@
    Diffusion: obtain $\mathbf{U}^{n+\frac{1}{2}}$ from heat equation with forcing term $\mathbf{U}^{n+\frac{1}{2}}=\mathbf{U}^{n}-\text{dt}{\mathbf{L}\mathbf{U}^{n+\frac{1}{2}}}+\mu\cdot(\mathbf{U}^{n}-\mathbf{U}_{\text{labeled}})$, applied $N_{s}$ times. We set $N_{s}=3$.
@@
    Projection to simplex: obtain $\mathbf{U}^{n+1}$ by projecting each row of $\mathbf{U}^{n+\frac{1}{2}}$ onto (3).
@@
    Displacement: Each row of the projected matrix from step 2 is replaced by its assigned vertex $\mathbf{e}_{k}$ in the simplex (3) depending on the decision threshold. This procedure is executed for every threshold in the given list.

### v3

@@ Abstract
Data sets with imbalanced class sizes, where one class size is much smaller than that of others, occur exceedingly often in many applications, including those with biological foundations, such as disease diagnosis and drug discovery. Therefore, it is extremely important to be able to identify data elements of classes of various sizes, as a failure to do so can result in heavy costs. Nonetheless, many data classification procedures do not perform well on imbalanced data sets as they often fail to detect elements belonging to underrepresented classes. In this work, we propose the BTDT-MBO algorithm, incorporating Merriman-Bence-Osher (MBO) approaches and a bidirectional transformer, as well as distance correlation and decision threshold adjustments, for data classification tasks on highly imbalanced molecular data sets, where the sizes of the classes vary greatly. The proposed technique not only integrates adjustments in the classification threshold for the MBO algorithm in order to help deal with the class imbalance, but also uses a bidirectional transformer procedure based on an attention mechanism for self-supervised learning. In addition, the model implements distance correlation as a weight function for the similarity graph-based framework on which the adjusted MBO algorithm operates. The proposed method is validated using six molecular data sets and compared to other related techniques. The computational experiments show that the proposed technique is superior to competing approaches even in the case of a high class imbalance ratio.
Keywords: Imbalanced data; molecular data; transformer; graph-based; data classification.
@@ 1 Introduction
In data classification, class imbalanced data is distinguished by an underrepresentation of one class (the “minority” class) in the data compared to the other class or classes (the “ majority” class(es)). As a result, data classification algorithms trained on imbalanced data tend to become biased with regard to the majority class(es) and often misclassify data points from the minority group, which is frequently the class of most interest 1, 2. This scenario is quite common in classification tasks, and it can be especially problematic if the minority class represents an important or critical outcome, such as in drug discovery, disease diagnosis, fraud detection, energy management, and other applications involving rare events 2. For example, drug discovery is a very exacting problem in science. In drug screening, the number of inactive molecules is typically thousands of times larger than that of active molecules. A successful drug must be highly potent, be free of serious side effects or off-target binding, and exhibit minimal or no toxicity 1. As a result, the majority of drug candidates are either inactive or unqualified, leading to highly imbalanced data sets 3, 4. Specifically, imbalanced data occur during virtual screening: identifying potential drug candidates from large compound libraries. Effective handling of imbalanced data can improve the identification of active compounds 5, 6, 7, 8. Another situation in which we deal with imbalanced data is quantitative structure-activity relationship modeling, which predicts a compound’s activity using the chemical composition 9, 10. Balancing the data can lead to more reliable predictions 11.
@@
Therefore, addressing imbalanced data is crucial for effective machine learning in many applications such as drug design and discovery. By employing a fusion of algorithm-level and data-level strategies, and focusing on appropriate evaluation metrics, researchers can build models that better identify rare but valuable compounds, leading to more successful outcomes in drug discovery and development.
@@
Techniques to deal with imbalanced data can be categorized into: 2, 12, 13, 14: data preprocessing or resampling methods15, 16, model-level techniques 2, 17, ensemble algorithms 18, and cost-sensitive methods 19. Strategies in the first group include resampling methods, which attempt to make the training set’s class distribution more balanced using sampling techniques, and feature selection methods, which are carried out in the feature space rather than the sample space 2. Examples of resampling procedures include oversampling 20, 21, undersampling 22, 23 and hybrid-sampling algorithms 15, 24, 25. Respectively, these procedures attempt to rebalance the sample space by creating new samples from the minority group, discarding representatives from the majority group, or a combination of both. Methods belonging to the second group, the model-level techniques, attempt to make adjustments to steps of the algorithm itself so that there is less bias towards classes of larger size. Ensemble algorithms use the output of multiple classifiers to make predictions, which can improve the performance of individual classifiers 2, 26. Lastly, cost-sensitive procedures instead assign a larger misclassification cost to classes of smaller size 19.
@@
Some cost-sensitive approaches to imbalanced data classification involve thresholding, which is applied in a postprocessing step 27. In particular, thresholding algorithms are cost-sensitive techniques which assign a probability to each of the test elements. For a threshold of 0.5, which is the default probability threshold value in most classifiers 28, any test element with a probability of at least 0.5 will be assigned to one class; otherwise, it will be assigned to another class. In the case of highly imbalanced data classification, the threshold of 0.5 may not be optimal; thus, threshold algorithms attempt to find a more optimal threshold boundary for the machine learning task. Usually, a balanced accuracy metric is used; some examples include the ROC curve (receiver operating characteristic curve) 29, the G-Mean 30, F1-score 28, the Matthews correlation coefficient 31 and the balanced accuracy 5. In these cases, the goal is to maximize the balanced accuracy metric. Overall, one approach for image classification and feature extraction is detailed in 32, which proposes a convolution-based efficient transformer image feature extraction network. Moreover, Wang et al. 33 introduce a micro-directional propagation model using deep learning, and Zhu et al. 34 focus on an adaptive agent decision model derived using autonomous learning deep reinforcement learning. In addition, Ref. 35 details feature extraction modules, while Ref. 36 presents a systematic review of picture fuzzy decision-making methods. Applications to sideband harmonics, internet market design, large-scale sliding puzzles are found in 37, 38, and 39.
@@
However, most of the cost-sensitive techniques for imbalanced data classification that search for the optimal threshold can be computationally extensive because retraining of the classifier is often required. Inspired by the work of Esposito et al. 1, in this paper, we detail an algorithm for imbalanced data classification which does not require any retraining in its procedure. In order to motivate the proposed procedure, we turn to the authors’ previous work 40, which demonstrated the success of their BT-MBO method for molecular data classification with very small labeled sets; refer to 41 for a survey of machine learning techniques for data with limited labeled elements in molecular science. The BT-MBO algorithm, which admits two dimensional molecular data in a special format called the simplified molecular input line entry specification (SMILES) 42 format, is composed of a bidirectional transformer 43 and a graph-based and modified Merriman-Bence-Osher (MBO) algorithm 44, 45 in series, utilizing the transformer model to generate molecular fingerprints which then serve as features for the MBO algorithm. The MBO method used in the BT-MBO model 40 generates a probability distribution over all classes for each data point, and it subsequently classifies each data element by choosing the class for which that element has the highest probability. However, in the case of imbalanced data, this procedure may cause the technique to miscategorize data elements in the minority group, since the probability distribution may be biased toward the majority classes. Thus, in the present work, we propose the BTDT-MBO method (bidirectional transformer with MBO techniques using a varying decision threshold), which is a novel adaptation of the BT-MBO technique for use on highly imbalanced molecular data.
@@
In particular, the proposed BTDT-MBO method incorporates a decision threshold adjustment in the MBO algorithm, which utilizes the probability distribution generated by the algorithm for each data point. Instead of following the typical MBO classification procedure, the BTDT-MBO algorithm tests a set of thresholds and chooses the threshold that produces the highest ROC-AUC score (area under the ROC curve) for the unlabeled points. The new thresholding step introduced in the BTDT-MBO method does not require any additional iterations of the MBO algorithm; rather, at the end of each iteration, we repeat the new thresholding step for each threshold. We refer to the modified MBO algorithm with decision threshold adjustment as the DT-MBO method. Another adaptation introduced in the BTDT-MBO method is the inclusion of the distance correlation 46, 47 as a weight function in the DT-MBO algorithm. The default weight function used in our BTDT-MBO method is a Gaussian kernel; in addition to experiments using a Gaussian weight function, we also carried out experiments using a distance correlation weight function.
@@
Given our previous BT-MBO method’s predictive success for the challenging scenario of scarcely labeled data 40, we focus on the difficult task of data classification of highly imbalanced molecular data sets in the present paper. Specifically, all six data sets used for benchmarking the proposed BTDT-MBO algorithm have an imbalance ratio (IR) of 16.5 or greater. We compare our results using BTDT-MBO to those of the GHOST procedure 1 applied to four common classification algorithms: random forest, gradient boosting decision trees, extreme gradient boosting, and logistic regression. Specifically, the GHOST technique 1 is an automated procedure that can be used in any data classification procedure and that adjusts the decision threshold of a classifier using the training set’s prediction probabilities (i.e., without needing to retrain the classifier); it achieved success in classifying imbalanced molecular data, demonstrated on data with varying levels of class imbalance. Detailed results on the six highly imbalanced data sets, as well as a discussion, are given in Section 3. The computational experiments indicate that our BTDT-MBO method obtains superior results to those of the comparison techniques.
@@
We present the main contributions of this work:
  • We present a new technique, the BTDT-MBO method, for molecular data classification, especially designed to be very useful even in cases of highly imbalanced data sets. This procedure incorporates Merriman-Bence-Osher techniques, as well as bidirectional transformers, distance correlation, and adjustments in the classification threshold.
  • The proposed BTDT-MBO algorithm is able to perform extremely well even in the cases of molecular data that is highly imbalanced and where the class sizes vary significantly. This is extremely important, since highly imbalanced data sets are often found in practice.
  • The proposed method utilizes several techniques to achieve good performance when using data with high class imbalance ratios, including adjustments in the classification threshold.
  • The proposed technique is evaluated using experiments on six data sets. It is shown that the BTDT-MBO method performs more accurately than the comparison methods in almost all cases. The ROC-AUC score is used as a metric.
@@ 2 Methods
In this section, we derive the proposed techniques. In particular, in Section 2.1, we outline the details of the proposed BTDT-MBO method (bidirectional transformer with MBO techniques 44, 45 using a varying decision threshold), which adapts some of the techniques of our previous work 40. In Section 2.2, we further discuss the distance correlation 47, 46 as a potential similarity metric for use in the BTDT-MBO method.
@@ 2.1 BTDT-MBO Algorithm
The authors’ previous work 40 demonstrated the success of their proposed BT-MBO method in molecular classification problems with small sets of labeled nodes, with the BT-MBO procedure outperforming state-of-the-art techniques on some benchmark data sets when only one percent of all data elements are viewed as labeled. The present work proposes a method, called BTDT-MBO, that adapts some of the procedures of the BT-MBO procedure for the data classification task on molecular data sets where the classes of the data set vary vastly in size.
Inspired in part by the success of Esposito et al. 1 in adjusting the classification threshold for machine learning classifiers to make predictions about imbalanced molecular data, the proposed BTDT-MBO method utilizes the aspects of the BT-MBO algorithm from 40 along with a decision threshold adjustment at each iteration of the method. We denote the adjusted MBO method with varied decision threshold by DT-MBO. The proposed BTDT-MBO algorithm consists of a bidirectional transformer (created by Chen et al. 43) and a DT-MBO algorithm in series, where the fingerprints (BT-FPs) generated by the transformer serve as input features to the DT-MBO algorithm. Figure 1 illustrates the steps of the BTDT-MBO algorithm. The BTDT-MBO algorithm is further outlined in Algorithm 2, and details of the DT-MBO method are given in Algorithm 1.
Figure 1: Flowchart illustrating the BTDT-MBO method.
@@ 2.1.1 MBO Scheme with Decision Threshold Adjustment
Here, we establish the similarity graph-based structure our BTDT-MBO scheme uses, which adapts techniques outlined in the literature 48, 49, 44. First, consider a graph involving the vertex set $V$ representing the data set nodes. We can then define a weight function $w:V\times V\to\mathbb{R}$, whose value represents the degree of similarity between two vertices of the graph. Specifically, $w(i,j)$ measures the degree of similarity of data elements $i$ and $j$ of the data set. While there are many choices for the weight function, it should be constructed so that the value of $w(i,j)$ corresponds to the extent of resemblance or closeness of the data elements $i$ and $j$. A popular weight function is the Gaussian weight function:
@@
where $d(i,j)$ represents the distance between vertices $i$ and $j$ (computed using some prespecified metric such as the Euclidean distance), and $\sigma>0$ 44, 49.
In this paper, we utilize the Gaussian weight function for some of our BTDT-MBO experiments, and for other experiments, we use the distance correlation 46 as a weight function. The distance correlation is defined and further discussed in Section 2.2.
Given a weight function $w$, let the weight matrix $\mathbf{W}$ be the matrix $\mathbf{W}_{ij}=w(i,j)$. Then let the degree of a vertex $i\in V$ be defined as $d_{i}=\sum_{j\in V}w(i,j)$. If we let $\mathbf{D}$ be a diagonal matrix containing diagonal elements $\{d_{i}\}$, then we can define the graph Laplacian as
@@
For outlining the Merriman-Bence-Osher (MBO) framework, formerly introduced in 45 to generate approximations to mean curvature flow and then applied using a graph-based setting to various applications such as hyperspectral imaging in works such as 48, 44, 50, 51, 52, we consider a general data classification problem with $m$ classes. Overall, the MBO algorithm ultimately aims to compute a label matrix $\mathbf{U}=(\mathbf{u}_{1},\dots,\mathbf{u}_{N})^{T}$; a row $\mathbf{u}_{i}\in\mathbb{R}^{m}$ of the label matrix contains the probability values of the data element $i$ belonging to each of the $m$ classes. For instance, the first element of $\mathbf{u}_{i}$ represents the probability that node $i$ is a member of the first class, the second element of $\mathbf{u}_{i}$ represents the probability that node $i$ is a member of the second class, and so on. Therefore, each vector $\mathbf{u}_{i}$ can be viewed as a component of the $m^{th}$ Gibbs simplex:
$\Sigma^{m}:=\{(y_{1},...,y_{m})\in[0,1]^{m}\;\;\text{such that}\;\;\sum_{k=1}^{m}y_{k}=1\}\vspace{0.2cm}$ (3)
@@
The vertices of $m^{th}$ Gibbs simplex are unit vectors $\{\mathbf{e}_{i}\}$, where the probability of belonging to class $i$ is 1, while the probability of belonging to any other class is 0. Thus, the vertices $\{\mathbf{e}_{i}\}$ correspond to data elements that belong exclusively to each of the $m$ classes. In our data classification problem with labeled and unlabeled data elements, the labeled data elements are assigned to their corresponding vertices of the Gibbs simplex.
@@
The process by which the MBO algorithm generates the optimal label matrix $\mathbf{U}$ relies on energy minimization, a time-splitting scheme, and transference to a graph-based setting as described above. A general graph-based machine learning algorithm for data classification can be formulated as minimizing the energy consisting of the sum of a regularization term that incorporates weights in addition to a fidelity part that integrates the labeled nodes 49.
@@
As Garcia and coauthors demonstrated in 44, one successful choice for the regularization term R is the Ginzburg-Landau functional 53, 54, often utilized in image processing tasks and associated applications due to its relation to the total variation functional 48, 55. For the fidelity term, one can choose an $L_{2}$ fit to all labeled nodes. In the continuous setting, $L_{2}$ gradient descent can then be applied to minimize the energy. This results in a modified Allen-Cahn equation containing an additional forcing part. Next, a time-splitting scheme produces a technique that alternates between two steps: a diffusion step, which uses the heat equation with an addition term, and a thresholding step. Finally, one can use the techniques outlined in 44, 48, 56 to transfer the procedure to a graph-based setting, in which the thresholding step is replaced by projection to the Gibbs simplex (3), followed by displacing to the nearest vertex in the $m^{th}$ Gibbs simplex. Additionally, to adapt to the graph setting, one can replace the Laplace operator by the graph Laplacian (2) or some normalization of the graph Laplacian.
@@
To formulate our decision threshold technique, we are inspired by Esposito et al. 1, which demonstrated success in classifying imbalanced molecular data by developing a procedure that adjusts and optimizes the decision threshold of machine learning classifiers. For binary data, the default decision threshold is usually 0.5, where the model assigns a sample to the minority class if its probability of belonging to the minority class is greater than 0.5 (this is identical to assigning the sample to the class for which it has the highest probability). However, machine learning classifiers are likely to misclassify minority points since the probability distribution over the two classes for a given data point may be skewed toward the majority class 57, so Esposito et al. 1 tested potential decision thresholds from 0.05 to 0.5 with a spacing of 0.05. Similarly, in the present work, we augment the MBO method by a process that optimizes the decision threshold for the displacement step, resulting in our proposed decision-threshold-MBO (DT-MBO) algorithm.
@@
Overall, we summarize the main iterative steps of the DT-MBO procedure for the data classification task used in the present work below, with more details given in Algorithm 1. The process is adapted from the MBO algorithm used in our previous work 40, with an additional decision threshold step in the displacement step. The below steps require a given labeled/unlabeled partition of a data set as well as a list of thresholds to test.
@@
    Diffusion step: compute $\mathbf{U}^{n+\frac{1}{2}}$ using the heat equation with an additional term using
    $\mathbf{U}^{n+\frac{1}{2}}=\mathbf{U}^{n}-\text{dt}{\mathbf{L}\mathbf{U}^{n+\frac{1}{2}}}+\mu\cdot(\mathbf{U}^{n}-\mathbf{U}_{\text{labeled}}),$
    applied $N_{s}$ times. Here, $\text{dt}>0$, and $\mathbf{U}_{\text{labeled}}$ is a matrix where a row corresponding to a labeled element is an indicator vector $\mathbf{e}_{k}$ if the node is of class $k$, and a zero vector for non-labeled points. Moreover, $\mathbf{\mu}$ is a vector whose entries are equal to a constant for labeled points, and zero otherwise. We set $N_{s}=3$.
@@
    Projection step: compute $\mathbf{U}^{n+1}$ via projecting (each row) of the matrix $\mathbf{U}^{n+\frac{1}{2}}$ onto (3).
@@
    Displacement step: Replace every row of the result from step 2 by its assigned vertex $\mathbf{e}_{k}$ in the $m^{th}$ Gibbs simplex (3) depending on the decision threshold. This procedure is executed for every threshold in the list.